OpenStudy (calvinnivlac):

Please help me fan, and medal. Could I please have the answer no explanation needed Write an equation of a line perpendicular to y = 1 over 2x − 5 in slope-intercept form that passes through the point (2, −7). y= 1 over 2x − 7 y= 2x + 2 y= −2x − 5 y= −2x − 3

OpenStudy (calvinnivlac):

I don't know what to do please help

OpenStudy (calvinnivlac):

@TuringTest

OpenStudy (calvinnivlac):

@Michele_Laino

OpenStudy (calvinnivlac):

@perl

OpenStudy (michele_laino):

hint: two lines are perpendicular each other, when the subsequent relationship holds: \[\Large m \times m' = - 1\] where m and m' are the slopes of those lines. Now, we have m=1/2, so what is m'=...? being m' the slope of the line perpendicular to your line

OpenStudy (calvinnivlac):

I am highly confused and I am in a rush because swimming practice can I please have the answer

OpenStudy (michele_laino):

I'm sorry I can not give you the answer directly, since the Code of Conduct

OpenStudy (calvinnivlac):

not to be rude or anything but I have to go and after this I will have 3 more questions >_<

OpenStudy (calvinnivlac):

better hint?

OpenStudy (calvinnivlac):

I don't remember this in my course. but if I don't submit the assessment in 20 minutes it will automatically submit without me

OpenStudy (michele_laino):

hint: \[\frac{1}{2} \times m' = - 1\]

OpenStudy (calvinnivlac):

a?

OpenStudy (michele_laino):

I understand you, nevertheless you have to understand me too. I have to respect the Code of Conduct.

OpenStudy (calvinnivlac):

what is m' equal to?

OpenStudy (calvinnivlac):

it must be B

OpenStudy (michele_laino):

 if I multiply both sides by 2, I get that. What is m'=...?